a rowing team rowing with the current, traveled 40km in 2 hrs. rowing against the current the team can only travel 16 km in 2 hrs find the team rowing rate in calm water and the rate of the current
Answer by mananth(16946)

Rowing speed =x mph
current speed =y mph
against current 2 hours
with current 2 hours

Distance with current 40 miles distance against current 16
t=d/r against current (x-y)
16.00 / ( x - y )= 2.00

2.00 x - -2.00 y = 16.00 ....................1
with current (x+y)
40.00 / ( x + y )= 2.00
2.00 ( x + y ) = 40.00
2.00 x + 2.00 y = 40.00 ...............2
Multiply (1) by 1.00
Multiply (2) by 1.00
we get y
2.00 x + -2.00 y = 16.00
2.00 x + 2.00 = 40.00
4.00 x = 56.00
/ 4.00
x = 14 mph

plug value of x in (1) y
2 x -2 y = 16
28 -2 -28 = 16
-2 y = 16
-2 y = -12 mph
y = 6
Rowing 14 mph
current 6 mph